OpenStudy (hysusonic):

can someone help me understand how to get a system of equations from a graph?

OpenStudy (hysusonic):

the teacher won't explain it and the lesson doesn't help either

OpenStudy (luigi0210):

Do you mean like a solution to them? Or just how to find an equation for each?

OpenStudy (hysusonic):

just how to find them

OpenStudy (phi):

Hopefully, the graph shows *straight lines* does it ?

OpenStudy (hysusonic):

 something similar to this

OpenStudy (phi):

you need to find the equation of each line. to find the equation of a line, the easiest way is to find any two points on the line. can you read off a point on one of the lines ?

OpenStudy (hysusonic):

i also have the curved line ones but i'll save that for later

OpenStudy (hysusonic):

the only point is (-1,-4)

OpenStudy (phi):

lines have lots of points

OpenStudy (hysusonic):

oh that's what you meant

OpenStudy (hysusonic):

that is the graph but i just want to learn how to actually find them for later

OpenStudy (phi):

ok (-1,-4) is on both lines (which means it is the solution) for the lower line, can you figure out what it's "y" value is when x is 3 ?

OpenStudy (hysusonic):

(3,0)?

OpenStudy (phi):

yes now we have two points, (3,0) and (-1, -4)

OpenStudy (hysusonic):

so i find the slope?

OpenStudy (phi):

we now find the slope between those two points

OpenStudy (hysusonic):

\[\frac{ -1-3 }{ -4-0 }\]

OpenStudy (phi):

yes, and that simplifies to -4/-4 or just 1

OpenStudy (hysusonic):

1

OpenStudy (hysusonic):

oh you meant it simplifies to both

OpenStudy (phi):

so the equation is y= mx + b we found m=1, so y = 1*x + b or (not bothering to show multiply by 1): y= x+b to find b, use (3,0) (or -1,-4)

OpenStudy (hysusonic):

wouldn't "b" be 3? because it's the y-intercept? or am i wrong

OpenStudy (phi):

we can read off the y-intercept from the graph, but it is not 3

OpenStudy (hysusonic):

oh wait y-intercept is when x=0 sorry brain fart

OpenStudy (phi):

y-intercept means where the line "intercepts" (fancy word for meets) the y-axis y-axis is the vertical axis

OpenStudy (phi):

using algebra and point (3,0) 0 = 3 + b add -3 to both sides -3+0 = 3+-3 + b -3 = b so b is -3 (and that is what the graph shows) y= x - 3 now put this in standard form: x - y = 3

OpenStudy (hysusonic):

-y=-x+3 and then divide the whole thing by -1?

OpenStudy (phi):

yes you can divide by -1 but standard form is A x + B y = C where A is positive so I would add +x to both sides , to change -y = -x +3 to x - y = 3 that is probably the form they want for the equation (variables in alphabetical order, first coefficient positive)

OpenStudy (phi):

now find the equation for the steep line. looking at the graph we see its "b" (y-intercept) is 3 so you just need the slope.

OpenStudy (phi):

you can see the line moves "over 1" , and up 7 so its slope will be 7 (change in y divided by change in x)

OpenStudy (hysusonic):

okay i'm a little bit lost.... what do i do after i put it in standard form?

OpenStudy (hysusonic):

oh wait never mind i see what you mean

OpenStudy (hysusonic):

so the answer to this one would be 7x+3=x-3?

OpenStudy (phi):

the second line will have an equation y= 7x + 3 and in standard form 7x -y = -3 and the system of equations is x - y = 3 7x -y = -3

OpenStudy (phi):

to solve, we can subtract the two equations x - y = 3 -(7x -y = -3) x-7x -y +y = 3 - -3 -6x = 6 x= 6/-6 x= -1 (as we know it would be) to find y, use either equation, with x= -1 example: x - y = 3 -1 - y= 3 -y = 4 y = -4

OpenStudy (hysusonic):

okay i see that but how do i get an answer like x+4=x+5?

OpenStudy (phi):

I think you are asking, How to solve the system of equations One way is elimination (see my post up above) the other way is substitution if we write both equations in slope-intercept form: y= x - 3 y = 7x + 3 the x-3 equals y and 7x+3 equals y so (where the lines meet and both lines have the same y value) x-3 = 7x +3 you can solve for x by adding -x to both sides x-x -3 = 7x -x + 3 -3 = 6x + 3 and you can finish solving ...

OpenStudy (hysusonic):

these are the answer choices they gave us

OpenStudy (phi):

ok, then they must want x-3 = 7x +3 or 7x + 3 = x - 3

OpenStudy (hysusonic):

so i was right?

OpenStudy (phi):

yes, up above you picked that out. notice we could have "cheated" by looking at the graph and seeing the solution is at x= -1 (and we expect to find y= -4) so if we put x=-1 into each choice, only one of them will show -4 = -4 after we simplify

OpenStudy (hysusonic):

so trial and error?

OpenStudy (phi):

depends. If you get stuck on a question or are lost, then testing each choice is a different way to solve the problem.

OpenStudy (hysusonic):

okay so the first way is to find the slope and y-intercept and the second way is plugging in the numbers

OpenStudy (phi):

yes

OpenStudy (hysusonic):

thank you phi.... can i bother you for one more type of graph?